In a group of 50 students, 40 students study Math, 32 study Physics and each student studies at least one of the subjects. Use this information to answer the questions below


2d) If a student is randomly selected, what is the probability that the student studies Math given that he or she DOES NOT study Physics?





1
Expert's answer
2021-06-13T15:44:25-0400
"N(A\cup M)=50"

"N(A)=40, N(B)=32"

"N(A\cap B)=N(A)+N(B)-N(A\cup B)"

"=40+32-50=22"

"N(A\cap B')=40-22=18"

"N(B\cap A')=32-22=10"


"P(B|A')=1"

The probability that the student studies Math given that he or she DOES NOT study Physics is 1.